OpenStudy (anonymous):

Then sum of two numbers is 50. The 2nd number is 5 times the 1st. find the number??

OpenStudy (anonymous):

x+y=50 y=5x

OpenStudy (anonymous):

so x=8 and 1/3 or \[\frac{25}{3}\]

OpenStudy (anonymous):

then put that into the equations and solve for y (you'll have to change 50 into thirds) or you can just multipy that number by five to see that it is \[y=\frac{125}{3}\] adding them together you get \[\frac{25}{3}+\frac{125}{3}=\frac{150}{3}=50\]

OpenStudy (anonymous):

always check at the end though... that's how you find out whether you mssed up

OpenStudy (anonymous):

Don't get it mate

OpenStudy (anonymous):

I've written down on my paper 5x+x=50

OpenStudy (anonymous):

thats correct

OpenStudy (anonymous):

now solve for x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

25/3?

OpenStudy (anonymous):

oh how I see get what you have written